OpenStudy (anonymous):

Finds the height of a triangle with a base of 35 ft and an area of 262.5 ft2

OpenStudy (anonymous):

We can use the area of a triangle formula. Formula: \(V = \dfrac{1}{2} b h\) Plug in what we know: \(262.5= \dfrac{1}{2} (35) h\) Multiply 1/2 * 35: \(262.5 = 17.5h\) Now divide 17.5 to both sides, what's 262.5 / 17.5? @aljonaeeee

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

Yep! So the height of the triangle is 15ft.

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @aljonaeeee